# Dynamic watermarks in images generated by diffusion models

### v1

@@ Abstract
High-fidelity text-to-image diffusion models have revolutionized visual content generation, but their widespread use raises significant ethical concerns, including intellectual property protection and the misuse of synthetic media. To address these challenges, we propose a novel multi-stage watermarking framework for diffusion models, designed to establish copyright and trace generated images back to their source. Our multi-stage watermarking technique involves embedding: (i) a fixed watermark that is localized in the diffusion model’s learned noise distribution and, (ii) a human-imperceptible, dynamic watermark in generates images, leveraging a fine-tuned decoder. By leveraging the Structural Similarity Index Measure (SSIM) and cosine similarity, we adapt the watermark’s shape and color to the generated content while maintaining robustness. We demonstrate that our method enables reliable source verification through watermark classification, even when the dynamic watermark is adjusted for content-specific variations. Source model verification is enabled through watermark classification. o support further research, we generate a dataset of watermarked images and introduce a methodology to evaluate the statistical impact of watermarking on generated content.Additionally, we rigorously test our framework against various attack scenarios, demonstrating its robustness and minimal impact on image quality. Our work advances the field of AI-generated content security by providing a scalable solution for model ownership verification and misuse prevention.
@@ 1 Introduction
Driven by increasing demands, deep generative models have gained widespread attention in academic, industry and public domains. Recent advances have made photorealistic image generation more accessible (Bińkowski et al., 2018; Heusel et al., 2017; Salimans et al., 2016; Zhou et al., 2019) as exemplified by DALL·E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022) and FLUX models (Labs, 2024). These models have also spurred the development of numerous image editing tools and text-to-video models including ControlNet (Zhang & Agrawala, 2023), Instruct-Pix2Pix (Brooks et al., 2022) and SORA (Brooks et al., 2024). While these models represent valuable digital assets, their widespread use raises critical ethical concerns, such as the potential for misuse in spreading false narratives, generating harmful representations, or infringing intellectual property (IP) rights.
To address these challenges, blind watermarking has emerged as a key strategy for IP protection and misuse prevention. Existing methods can be broadly categorized into two types: (1) embedding watermarks directly into the generative model to enable source verification (Wen et al., 2023; Fernandez et al., 2023; Peng et al., 2025; Min et al., 2024), and (2) embedding watermarks into generated images to protect content from unauthorized use (Zhao et al., 2023; Fernandez et al., 2023; Liu et al., 2023). However, different generation models and platforms may use varying content creation techniques. The diversity of generative models and platforms, each employing distinct algorithms, training processes, and representation formats, complicates the development of a universal watermarking solution. In addition, current watermarking methods face limitations. Moreover, fixed watermarks embedded during training are often predictable and vulnerable to removal or tampering through reverse engineering or image-processing techniques (Moffat, 2019; Bińkowski et al., 2018; Heusel et al., 2017; Salimans et al., 2016; Zhou et al., 2019).
Figure 1: Overview of the proposed dual watermarking framework. The first stage embeds a fixed QR-code watermark directly into the diffusion model’s latent space, ensuring model ownership verification. The second stage introduces a dynamic watermarking mechanism that imperceptibly modifies generated images, enhancing traceability while maintaining image quality.
We propose an intuitive dual watermarking strategy, embedding: (i) a unique, model-specific QR code watermark directly into the diffusion model and, (ii) dynamic watermarks into images generated by the watermarked diffusion model. Our fixed QR-code watermark uniquely encodes model-specific metadata, within the diffusion model, including the IP address of machine it was trained on and date-time information. This ensures uniqueness, resistance to tampering, and complements dynamic watermarks in generated images for robust traceability through mutual verification. Our dynamic watermarking process that adjusts within feature and pixel spaces, thereby enhancing robustness through dynamic transformations based on the generated content. In the feature space, The high-level feature representations are extracted from the original and watermarked images using a pre-trained feature extractor. We calculate the cosine similarity (Rahutomo et al., 2012) between the features extracted from the images before and after applying watermark transformations. This step aims to enhance the separation of images by analyzing their feature similarities. In the pixel space, we evaluate the quality of images using the Structural Similarity Index (SSIM) (Channappayya et al., 2008), which quantifies the perceived differences between the original and transformed images. By mitigating the structural impact of the watermark embedding, we ensure that generated images are still properly classified post-watermarking.
Moreover, extensive research has shown that edges, frequency, and texture are the three key elements that influence human perception of subtle differences in images (Afchar et al., 2018; Chen et al., 2022; Cozzolino et al., 2017; Do & Vetterli, 2005; Chen et al., 2024). Leveraging this insight, we introduce a novel validation method to evaluate the impact of blind watermarking on image quality, utilizing 11 image statistics for assessment. The key contributions of this paper include:
  • A dynamic watermarking strategy that enhances IP protection for generative models and enables traceability of generated content.
  • An intuitive embedding methodology based on SSIM and cosine similarity, dynamically adjusting watermarks in feature and pixel spaces to improve robustness and minimize visual impact.
  • A fixed QR-code watermark embedded within the diffusion model, cross-verifiable with dynamic watermarks in generated images, ensuring tamper resistance and content provenance.
  • A novel blind watermark impact assessment mechanism that evaluates watermark visibility using 11 image statistics, addressing human-perceptible factors such as edges, frequency, and texture.
By addressing the limitations of existing methods and providing a robust, scalable solution, our work advances the field of AI-generated content security and contributes to ongoing efforts to mitigate ethical concerns in generative AI.
@@ 2 Related Work
Fundamentally, diffusion models are generative frameworks built on the diffusion principles observed in non-equilibrium thermodynamics. Generally, these are implemented through a finite Markov chain (Geyer, 1992) for forward and reverse diffusion. Recently, diffusion models have been adapted for conditional image generation tasks, such as inpainting and text-guided generation and editing. Their iterative de-noising steps enables zero-shot image editing by guiding the generative process (Heusel et al., 2017; Salimans et al., 2016; Zhou et al., 2019).Watermarking techniques are actively researched to address copyright and misuse concerns in diffusion-based generative models. Here, we focus on the two main approaches, i.e, embedding watermarks within the model and, embedding watermarks in generated images.
Early research proposed methods for watermarking GANs by constructing mappings between trigger inputs and generator outputs with regularization constraints (Yu et al., 2021; Ong et al., 2021; Fei et al., 2022). However, these techniques cannot be directly applied to diffusion models due to architectural differences (Creswell et al., 2018; Cao et al., 2022). Additionally, Wen et al. proposed a robust tree-ring watermarking method against multiple image transformations and attacks (Wen et al., 2023). In addition, Li et al. introduced a diffusion-based watermarking approach, DiffWA (Li, 2023), which leverages the diffusion process to embed noise- and tamper-resistant watermarks.
In generative models, many works attempt to embed watermarks within the training dataset to protect the generative model’s training data (Chai et al., 2020). However, this approach may be inefficient, as embedding new information requires costly retraining. Some research combines watermark embedding with the generative process  (Gragnaniello et al., 2021; Wang et al., 2020), aligning more closely with model watermarking techniques. However, this approach faces two key limitations: (i) it applies primarily to GANs, while latent diffusion models (LDMs) are increasingly replacing GANs for most applications and, (ii) the watermark is embedded from the start of model training (Fei et al., 2022; Lin et al., 2022), a strategy that is challenging to sustain, given the resource-intensive nature of generative model training. Studies have shown that time-optimized fine-tuning of a generative model’s latent decoder combined with an appropriate watermark extractor can achieve effective watermarking results (Fernandez et al., 2023).
@@ 3 Proposed Method
In this work, we propose two watermark embedding branches. (i) A model watermarking branch that embeds a fixed QR-code watermark into the learned noise distribution of the diffusion model, ensuring robust traceability and ownership verification. (ii) An image watermarking branch that embeds dynamic watermarks into generated images, balancing imperceptibility and resilience to attacks. Together, these branches enable model ownership verification and misuse prevention.
@@ 3.1 Model Watermarking
Figure 2: To verify model ownership, we embed a QR-code watermark into a target diffusion model, leveraging a watermark retrieval key to isolate watermark data and image data distributions. (Left) Through combined losses, the training stage contains both watermark and image diffusion stages to construct a common, watermarked latent space. (Right) Sampling from the watermarked model will generate images as usual. When provided with the watermark retrieval key as an input in the sampling process, the watermark will be extracted, allowing for model identification.
We employ the Stable Diffusion Model (SDM) (Rombach et al., 2022) for our experiments. To enable watermark embedding, we modify the latent diffusion process by introducing a modified Gaussian kernel. The data distribution is defined as $q(z)$, where $z$ represents data in the latent space.
$q(z_{t}|z_{t-1})=N\left(z_{t};\sqrt{\alpha_{t}}(z_{t-1}+\phi_{t}),\eta^{2}(1-\alpha_{t})I\right),$ (1)
Here, $\alpha_{t}$ determines the step size for each diffusion step. $\phi_{t}$ is a constant schedule parameter in the diffusion process, which adjusts the input signal at each timestep $t$. The scheduling function can shift the input data, to assist in extracting and representing input features at specific timesteps. $\eta\in[0,1]$ controls the noise scale and is a constant within the range $[0,1]$.
The Watermark Diffusion process serves as an extension of the traditional SDM diffusion process, introducing a trigger key, which we denote as ‘$\kappa$’ (see Fig. 2) to alter the diffusion pathway for the variable $z_{t}$ such that:
@@
where $\gamma_{\kappa}$ is a blending factor that modulates the influence of the watermark on the generated output. The reverse diffusion process retrieves the embedded watermark from the latent space for verification. The trigger key is integrated during sampling, as shown in Fig. 2.
Watermark Embedding can be accomplished by fine-tuning the host model $\epsilon_{\theta}^{o}$. In each iteration, we sample a data instance $z_{0}$ from the training dataset $D_{\text{train}}$ and a watermark example $z_{0}^{\text{w}}$ from the watermark dataset $D_{\text{wm}}$. Noise samples $\epsilon$ and $\epsilon_{\text{w}}$ are then drawn from $N(0,I)$ separately for task and watermark data, along with a timestep $t$ sampled from $\text{Uniform}(\{1,...,T\})$.
For the sample $z_{0}$, the latent representation $z_{t}$ at timestep $t$ is computed using the latent diffusion process. For the watermark sample $z_{0}^{\text{w}}$, we first compute its latent representation $z_{t}^{\text{w}}$ within the latent diffusion process, and then construct the state $\tilde{z}_{t}^{\text{w}}$ in the Watermark Diffusion Process based on $z_{t}^{\text{w}}$ as follows:
@@
$\begin{split}L_{WDP}={E}_{t\sim[1,T],z_{0},z_{w0},\epsilon_{t}}\Big{[}&\,\gamma_{\epsilon}\|\epsilon-\epsilon_{\theta}(z_{t},t)\|^{2}+\|\epsilon_{w}-\epsilon_{\theta}(\hat{z}_{wt},t)\|^{2}\Big{]},\end{split}$ (4)
Here, $\gamma_{\epsilon}$ is a weighting factor that balances the standard diffusion process $\epsilon$ and the watermark diffusion process $\epsilon_{w}$ during training.
Watermark Extraction to verify model ownership, can be achieved using the standard reverse diffusion process given the common latent space that was constructed through the combined diffusion processes discussed previously. The resulting output should be a reconstruction of the QR-code watermark that was used in the joint-training process.Given a model $\epsilon_{\theta}$, along with the known trigger $\kappa$ and trigger factor $\gamma_{\kappa}$, we first sample $z_{T}^{\text{w}}$ from $N(0,1)$. Next, we calculate its corresponding state $\tilde{z}_{t}^{\text{w}}$ in the watermark diffusion process as follows:
@@
$\begin{split}{z}_{t-1}^{\text{w}}=\frac{1}{\sqrt{\alpha_{t}}}\left[{z}_{t}^{\text{w}}-\frac{(1-\alpha_{t})}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\gamma_{\kappa}\hat{z}_{t}^{\text{w}}+(1-\gamma_{\kappa})\kappa,t)\right]+\sigma_{t}z,\end{split}$ (6)
where $\sigma_{t}$ is the noise level at timestep $t$ and $z$ is a random sample from $N(0,I)$. The selection of the trigger $key$ and the factor $\gamma_{\kappa}$ should ensure sufficient divergence between the state distribution in the Watermark Diffusion Process and that in the standard diffusion process.
Figure 3: Overview of the dynamic watermarking framework, structured in three stages: (A) Training Stage: Fine-tunes the SDM decoder with a pre-trained watermark extractor to embed dynamic watermarks. (B) Extractor Training Stage: Trains the watermark extractor to encode and retrieve binary watermark information, maintaining reliability under transformations. (C) Inference Stage: Generates watermarked images using the SDM decoder, allowing the extractor to recover embedded watermarks for verification.
@@ 3.2 Image Watermarking
A dynamic watermarking mechanism embeds a unique watermark for each image. Watermarks vary in shape and colour while preserving key features of the reference watermark. In the feature space, we compute the cosine similarity between the original and transformed watermark images. Minimising this similarity enhances orthogonality, making extraction or tampering more difficult. In the pixel space, we maximise the SSIM to ensure high visual fidelity. These strategies enhance watermark robustness, imperceptibility, and classifiability without compromising the generative capabilities. The cosine similarity between the original $f_{o}$ and transformed $f_{t}$ watermark feature vectors is defined as:
$\cos(\theta)=\frac{f_{o}\cdot f_{t}}{\|f_{o}\|\|f_{t}\|}.$ (7)
By minimizing the cosine similarity, we aim to make $f_{o}$ and $f_{t}$ as orthogonal as possible within the high-dimensional feature space. This optimization ensures that the transformed watermark maintains its semantic structure while being imperceptible in the generated image.
In the pixel space, we calculate the SSIM between the original and transformed watermark images, denoted as $I_{o}$ and $I_{t}$. The SSIM is defined as:
$\text{SSIM}(I_{o},I_{t})=\frac{(2\mu_{I_{o}}\mu_{I_{t}}+C_{1})(2\sigma_{{I_{o}}{I_{t}}}+C_{2})}{(\mu_{I_{o}}^{2}+\mu_{I_{t}}^{2}+C_{1})(\sigma_{I_{o}}^{2}+\sigma_{I_{t}}^{2}+C_{2})},$ (8)
where $\mu_{I_{o}}$ and $\mu_{I_{t}}$ are the means of $I_{o}$ and $I_{t}$, $\sigma_{I_{o}}^{2}$ and $\sigma_{I_{t}}^{2}$ are the variances, and $\sigma_{{I_{o}}{I_{t}}}$ is the covariance between ${I_{o}}$ and ${I_{t}}$. Constants $C_{1}$ and $C_{2}$ are used for stability. By maximizing SSIM, we ensure that the transformed watermark image maintains high visual similarity to the original in the pixel space, thereby preserving visual quality.
To ensure that the watermark meets the requirements in both feature space and pixel space, we formulate an optimization problem, minimizing the following objective:
$\min_{\text{transform parameters}}\left(\lambda_{cosine}\cdot\cos\theta-\lambda_{SSIM}\cdot\text{SSIM}(I_{o},I_{t})\right),$ (9)
Where $\lambda_{cosine}$ and $\lambda_{SSIM}$ are weighting parameters that control the relative importance of cosine similarity and SSIM in the objective function. The goal is to reduce $\cos\theta$ between the original and transformed feature vectors, thereby increasing orthogonality. Concurrently, maximizing the SSIM between the original and transformed images in the pixel space preserves visual quality.
Image-to-binary sequences. In Fig.  3 (A), the watermark image is first converted into a binary sequence. We employ Huffman coding-based character encoding to reduce the size of binary sequence Gajjala et al. (2020). The average encoding length $L_{\text{avg}}$ can be represented as: =
$L_{\text{avg}}=\sum_{i=1}^{n}p_{i}l_{i},$ (10)
where $p_{i}$ is the probability of character $c_{i}$, $l_{i}$ is the encoding length for $c_{i}$, and $n$ is the total number of characters in the set.
Image Watermark Embedding. Watermark embedding consists of two steps. First, we train a watermark extractor. Then, we fine-tune the SDM decoder to embed specific watermark information in all generated images.
As shown in Fig. 3(B), the watermark extractor $W_{E}$ training builds on HiDDeN (Hsu & Wu, 1999). We embed binary watermark information by jointly optimizing the parameters of the watermark encoder and $W_{E}$. Ultimately, only $W_{E}$ is retained as the watermark extractor. The watermark encoder takes a training image $x_{o}$ and a binary watermark message $m$ as inputs, producing a residual image $x_{\delta}$. The watermarked image is obtained by scaling $x_{\delta}$ with a factor $\alpha$. In each optimization step, a transformation $T$ is randomly selected from a predefined set (e.g., cropping, compression) to generate the watermarked image. The extractor $W_{E}$ then recovers the binary watermark $m^{\prime}$ as:
@@
The loss for the binary watermark information is calculated as:
$L_{E}=-\sum_{i=1}^{k}\left[m_{i}\cdot\log\sigma(m^{\prime}_{i})+(1-m_{i})\cdot\log(1-\sigma(m^{\prime}_{i}))\right]$ (12)
In Fig.  3 Part (A), the SDM uses the latent vector $z$ decoded by decoder to produce a generated image. To support multiple watermarks, the decoder $D_{m}$ is extended to accept both the latent vector $z$ and the condition vector $e_{i}$, which specifies which watermark to embed.
A training image is encoded by the SDM encoder $E$, and then combined with the embedding vector $e_{i}$ (derived from the condition $i$) to control the watermark embedding as follows:
@@
$L_{m}=-\sum_{j=1}^{k}\Big{[}$ ${m_{i}}^{(j)}\cdot\log\sigma\left({m^{\prime}_{i}}^{(j)}\right)+\left(1-{m_{i}}^{(j)}\right)\cdot\log\left(1-\sigma\left({m^{\prime}_{i}}^{(j)}\right)\right)\Big{]}$

### v2

@@ Abstract
High-fidelity text-to-image diffusion models have revolutionized visual content generation, but their widespread use raises significant copyright concerns. We propose a novel multi-stage watermarking framework for diffusion models, designed to establish copyright and trace generated images back to their source. Our technique involves embedding: (i) a fixed watermark that is localized in the diffusion model’s learned noise distribution and, (ii) a human-imperceptible, dynamic watermark in generates images, leveraging a fine-tuned decoder. By leveraging the Structural Similarity Index Measure (SSIM) and cosine similarity, we adapt the watermark’s shape and color to the generated content while maintaining robustness. We demonstrate that our method enables reliable source verification through watermark classification, even when the dynamic watermark is adjusted for content-specific variations. Source model verification is enabled through watermark classification. We generate a dataset of watermarked images and introduce a methodology to evaluate the statistical impact of watermarking on generated content. Additionally, we rigorously test our framework against various attack scenarios, demonstrating its robustness and minimal impact on image quality.
@@ 1 Introduction
Recent deep generative models have made photorealistic image generation more accessible [2, 16, 25, 29], as exemplified by DALL·E 2 [23], Stable Diffusion [24], and FLUX models [18]. These models have also spurred the development of numerous image editing tools and text-to-video models, including ControlNet [28], Instruct-Pix2Pix [3], and SORA [4]. While these models represent valuable digital assets, their widespread use raises the potential for misuse in spreading false narratives, generating harmful representations, or infringing intellectual property (IP) rights.
To address these challenges, blind watermarking has emerged as a key strategy for IP protection and misuse prevention. However, current watermarking methods face limitations, as fixed watermarks embedded during training are often predictable and vulnerable to removal or tampering through reverse engineering or image-processing techniques [21, 2, 16, 25, 29]. We propose a novel dual watermarking strategy by embedding: (i) a unique, model-specific QR-code watermark directly into the diffusion model, and (ii) dynamic watermarks into images generated by the model. The fixed QR-code watermark uniquely encodes model-specific metadata, including the IP address of the training machine and timestamp information.
Our dynamic watermarking process enhances robustness by dynamically adjusting watermark transformations based on generated content within both feature and pixel spaces. In the feature space, high-level representations are extracted from original and watermarked images using a pre-trained feature extractor. We then calculate cosine similarity [22] between these features to quantify their semantic consistency. In the pixel space, image quality is evaluated using the Structural Similarity Index (SSIM) [1, 6, 9, 10, 7]. We propose a validation method that assesses the impact of blind watermarking on image quality through an analysis of 11 distinct image statistics.
@@ 2 Related Work
Figure 1: We embed a QR-code watermark into a target diffusion model, leveraging a watermark retrieval key to isolate watermark data and image data distributions. (Left) The training stage contains both watermark and image diffusion stages to construct a common, watermarked latent space. (Right) Sampling from the watermarked model will generate images as usual. When provided with the watermark retrieval key as an input in the sampling process, the watermark will be extracted, allowing for model identification
Fundamentally, diffusion models are generative frameworks inspired by diffusion processes in non-equilibrium thermodynamics. Typically, these models implement forward and reverse diffusion processes via finite Markov chains [14]. Recently, diffusion models have been adapted for conditional image generation tasks, such as image inpainting, text-guided generation, and editing. Their iterative denoising steps enable zero-shot image editing by guiding the generative process [16, 25, 29].
In generative models, watermark embedding within training datasets has been explored to protect the model’s data [5]. However, this method can be inefficient, as adding new watermark information typically requires costly retraining. Alternative approaches integrate watermark embedding directly into the generative process [15, 26], closely aligning with model watermarking techniques. Yet, such methods have two primary limitations: (i) they predominantly target GAN-based models, despite latent diffusion models (LDMs) increasingly replacing GANs in practical applications; and (ii) watermarks are embedded during initial model training [11, 20], which is resource-intensive and difficult to maintain. Recent studies indicate that efficient watermark embedding can be achieved through optimized fine-tuning of the generative model’s latent decoder combined with an effective watermark extractor [12].
@@ 3 Proposed Method
We propose two watermark embedding branches: (i) a model watermarking branch that embeds a fixed QR-code watermark into the learned noise distribution of the diffusion model, ensuring robust traceability and ownership verification; and (ii) an image watermarking branch that embeds dynamic watermarks into generated images, balancing imperceptibility with resilience to attacks.
@@ 3.1 Model Watermarking
We employ the Stable Diffusion Model (SDM) [24] for our experiments. To enable watermark embedding, we modify the latent diffusion process by introducing a modified Gaussian kernel. The data distribution is defined as $q(z)$, where $z$ represents data in the latent space. The watermark diffusion process serves as an extension of the traditional SDM diffusion process, introducing a watermark retrieval key, which we denote as $\kappa$’ (see Fig. 1) to alter the diffusion pathway for the variable $z_{t}$ such that:
@@
where $\gamma_{\kappa}$ is a blending factor that modulates the influence of the watermark on the generated output. Watermark embedding can be accomplished by fine-tuning the host model $\epsilon_{\theta}^{o}$. In each iteration, we sample a data instance $z_{0}$ from the training dataset $D_{\text{train}}$ and a watermark example $z_{0}^{\text{w}}$ from the watermark dataset $D_{\text{wm}}$. Noise samples $\epsilon$ and $\epsilon_{\text{w}}$ are then drawn from $N(0,I)$ separately for task and watermark data, along with a timestep $t$ sampled from $\text{Uniform}(\{1,...,T\})$.
For the watermark sample $z_{0}^{\text{w}}$, we first compute its latent representation $z_{t}^{\text{w}}$ at timestep $t$ within the latent diffusion process, and then construct the state $\tilde{z}_{t}^{\text{w}}$ in the Watermark Diffusion Process based on $z_{t}^{\text{w}}$ as follows:
@@
$\begin{split}L_{WDP}={E}_{t\sim[1,T],z_{0},z_{w0},\epsilon_{t}}\Big{[}&\,\gamma_{\epsilon}\|\epsilon-\epsilon_{\theta}(z_{t},t)\|^{2}\\ &+\|\epsilon_{w}-\epsilon_{\theta}(\hat{z}_{wt},t)\|^{2}\Big{]},\end{split}$ (3)
Here, $\gamma_{\epsilon}$ is a weighting factor that balances the standard diffusion process $\epsilon$ and the watermark diffusion process $\epsilon_{w}$.
Figure 2: Overview of the dynamic watermarking framework, structured in three stages: (A) Training Stage: Fine-tunes the SDM decoder with a pre-trained watermark extractor to embed dynamic watermarks. (B) Extractor Training Stage: Trains the watermark extractor to encode and retrieve binary watermark information, maintaining reliability under transformations
Watermark extraction can be achieved using the standard reverse diffusion process given the common latent space that was constructed through the combined diffusion processes discussed previously. The resulting output should be a reconstruction of the QR-code watermark that was used in the joint-training process.Given a model $\epsilon_{\theta}$, along with the known trigger $\kappa$ and trigger factor $\gamma_{\kappa}$, we first sample $z_{T}^{\text{w}}$ from $N(0,1)$. Next, we calculate its corresponding state $\tilde{z}_{t}^{\text{w}}$ in the watermark diffusion process as follows:
@@
$\begin{split}{z}_{t-1}^{\text{w}}&=\frac{1}{\sqrt{\alpha_{t}}}\left[{z}_{t}^{\text{w}}-\frac{(1-\alpha_{t})}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\Bigl{(}\gamma_{\kappa}\hat{z}_{t}^{\text{w}}+(1-\gamma_{\kappa})\kappa,t\Bigr{)}\right]\\ &\quad+\sigma_{t}z,\end{split}$ (5)
where $\sigma_{t}$ is the noise level at timestep $t$ and $z$ is a random sample from $N(0,I)$. The selection of the trigger keykey and the factor $\gamma_{\kappa}$ should ensure sufficient divergence between the state distribution in the watermark diffusion process and that in the standard diffusion process.
@@ 3.2 Image Watermarking
Watermarks vary in shape and color while preserving key features of the reference watermark. In the feature space, we minimize the cosine similarity between original and transformed watermark features, enhancing orthogonality and robustness against extraction or tampering. Simultaneously, in the pixel space, we maximize the Structural Similarity Index Measure (SSIM) between original and transformed watermark images, ensuring high visual fidelity. These strategies enhance watermark robustness, imperceptibility, and classifiability without compromising generative capabilities.
To jointly satisfy constraints in both feature and pixel spaces, we formulate an optimization problem with the following objective:
$\min\left(\lambda_{cosine}\cdot\cos\theta-\lambda_{SSIM}\cdot\text{SSIM}(I_{o},I_{t})\right),$ (6)
where $\lambda_{cosine}$ and $\lambda_{SSIM}$ are weighting parameters that control the relative importance of cosine similarity and SSIM in the objective function. Original and transformed watermark images denoted as $I_{o}$ and $I_{t}$.
In Fig.  2 (A), the watermark image is first converted into a binary sequence. We employ Huffman coding-based character encoding to reduce the size of binary sequence [13]. Watermark embedding consists of two steps. First, we train a watermark extractor. Then, we fine-tune the SDM decoder to embed specific watermark information in all generated images.
As shown in Fig. 2(B), the watermark extractor $W_{E}$ training builds on HiDDeN [17]. We embed binary watermark information by jointly optimizing the parameters of the watermark encoder and $W_{E}$. Only $W_{E}$ is retained as the watermark extractor. The watermark encoder takes a training image $x_{o}$ and a binary watermark message $m$ as inputs, producing a residual image $x_{\delta}$. The watermarked image is obtained by scaling $x_{\delta}$ with a factor $\alpha$. The extractor $W_{E}$ then recovers the binary watermark $m^{\prime}$ as:
@@
In Fig. 2 Part (A), the SDM uses the latent vector decoded by decoder to produce a generated image. To support multiple watermarks, the decoder $D_{m}$ is extended to accept both the latent vector $z$ and the condition vector $e_{i}$, which specifies which watermark to embed. A training image is combined with the embedding vector $e_{i}$ (derived from the condition $i$) to control the watermark embedding as follows:
@@
$L_{m}=-\sum_{j=1}^{k}\Big{[}$ ${m_{i}}^{(j)}\cdot\log\sigma\left({m^{\prime}_{i}}^{(j)}\right)\quad+$
$\left(1-{m_{i}}^{(j)}\right)\cdot\log\left(1-\sigma\left({m^{\prime}_{i}}^{(j)}\right)\right)\Big{]}$ (9)